What are the steps for questions 2,3,4&5
2. The Dr. orders 1,000 Adrenaline (Epine mcg of Adrenaline (Epinephrine). How many ml should be prepared using a 30 mL vial labeled 1 mg in 1mL? ANSWER 3. Haldol (haloperidol) 2 mg is ordered. Using a TB syringe, how many mL should be prepared using an ampule labeled 5mg in 1 mL? ANSWER Shade the syringe with the proper dosage. 2 345 .678 9 1.0 4. The Dr orders didronel (Dtidronate) 300 mg q 6 hours. The medicine is available in a multidose vial labeled 50 mg in 1 mL. How many mL should be given per dose? ANSWER 5. Your patient is receiving Dycill (dicloxacillin Sodium) 1 gm po daily. Dycill is available in 500 mg capsules. How many capsules should be given? ANSWER
Answer #1

ANSWER

2. Calculation of amount of adrenaline prepared for use.

       * Formula

Amount of drug (ml) prepared = Desired dose(D) / Amount in hand (H) * Quantity(Q)

        Physician ordered dose(D) = 1000 mcg.

                                  1 mcg = 0.001 mg.

                              1000 mcg = 1 mg.

      Amount in hand(H)= 30 mg ( 1 mg in 1 ml)

                          Quantity(Q) = 30 ml.

As per formula,

        Amount of drug (ml) prepared = D / H * Q.

                                                              = 1 / 30 * 30

                                                          = 0.033 * 30

                                                              = 1 ml.

3. Calculation of amount of haloperidol prepared for use.

        * Formula

Amount of drug (ml) prepared = Desired dose / Amount in hand * Quantity.

         Desired dose (D) = 2 mg.

                      Amount in hand (H) = 5 mg.

                               Quantity (Q) = 1ml.

       As per formula,

           Amount of drug prepared = D / H * Q

                                                         = 2 / 5 * 1

                                                         = 0.4 * 1

                                                        = 0.4 ml.

4. Calculation of amount of didronel given per dose.

         * Formula.

Amount of drug given per dose (ml) = Desired dose / Amount in hand * Quantity.

                            Desired dose (D) = 300 mg (q 6hr)

                       Amount in hand (H) = 50 mg.

                            Quantity (Q) = 1 ml.

          As per formula,

             Amount of drug given per dose = D / H * Q

                                                                      = 300 / 50 *1

                                                                      = 6 * 1

                                                                      = 6 ml.

5. Calculation of number of capsules of Dycill needed for use.

       * Formula.

Number of capsules required = Desired dose / Amount in hand * Quantity

                     Desired dose(D) = 1 gm

                                           = 1000 mg.

                 Amount in hand(H) = 500 mg.

                        Quantity(Q) = 1 Capsule

       As per formula

         Number of capsules required = D / H * Q

                                                               = 1000 / 500 *1

                                                               = 2 * 1

                                                               = 2 capsules.